Want to make sure I got the correct answer! I got a=44(44,45,46)

Find three consecutive numbers such that the difference between the largest number and three-elevenths of the smallest number is thirty-four.

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

Let smallest =x
The next DISABLED_event_one= x+1
The last DISABLED_event_one= x+2
x+2 - (3/11) x =34 ;; subtract 3/11 from 1
(8/11) x =32
8x=352 ;; I can multiply by 11 or multiply by the reciprocal
x=44
numbers are 44, 45, and 46
(3/11)*44=12
46-12=34
Check

Answer by MathTherapy(10555)


Want to make sure I got the correct answer! I got a=44(44,45,46)

Find three consecutive numbers such that the difference between the largest number and three-elevenths of the smallest number is thirty-four.
Your answer is correct!! Great job!!